OpenStudy (anonymous):

I need help with my math homework. Translate each phrase and select the one that is DIFFERENT from the others. One-third of x The ratio of x and 3 The quotient of x and 3 The difference between x and 3

OpenStudy (anonymous):

one third of x = 1/3x the ratio of x and 3 = x/3 (which is the same as 1/3x the quotient of x and 3 = x/3 (also 1/3x) the difference between x and 3 = x - 3